It is known that 60% of riders on motor bikes wear helmets.The traffic police randomly looking for riders on motor bikes not wearing a helmet observe 5 riders on motor bikes.
a)Find the probability that he finds that every one of the riders was wearing a helmet.
b) Find the probability that at least two of the riders were wearing helmets.
c)What is the expectation of number of riders out of 5 riders wearing helmets?

Answer by edjones(8007)

p=.6, q=.4
(p+q)^5=p^5+5p^4q+10p^3q^2+10p^2q^3+5pq^4+q^5
.
a) p^5=.6^5=.07776 probability that he finds that every one of the riders was wearing a helmet.
.
b) 1-(5pq^4+q^5)=1-(.0768+.01024)=.91296 probability that at least two of the riders were wearing helmets.
.
c) 5*.6=3 expectation of number of riders out of 5 riders wearing helmets.
.
Ed